Created by Main.AntonioVergara on 2007-10-11 - 09:34
Sector 8-1
Going-on Activities
- SSS Replacement, including:
- Replacement and interconnection works
- Leak tests
- Local ElQA tests
- Inner Triplets:
- Left 1: closed, being leak tested
- Right 8: leak localised in C'. Repair going on. Final test wil be done once the SSS repairs are finished.
- Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.
- LHCb compensator tests will last till tomorrow Friday
- Installation of the instrumentation rods in DFBs Right 8 - COMPLETED
- Water pump check Point 8 planned for this week
Planned activities
- PIC-PC interface test - Wk 43
- Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48)
- ATLAS-RP cable installation - Wks 48-49
- DFBXH:
- Ineo cabling - Wk 43
- Installation of the removed water-cooled cable - Flohe Wk ?
- New cleaning of the DFBXH, the instrumentation will have to be removed and the DFBX re-validated by AT-MEL afterwards - Wk ?
- MB1158 diode: electrical continuity tests - Monday 15 Oct
- Flexible water cable reconnection DFBs Left 1 - Week 42
General
- AUG test on October 19 - Real impact will be described today
- Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped. The purging acitivities required to continue with he vacuum tests could be carried out using a temporary solution. Paul will ask Francois to go ahead with it.
- A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing
Comments:
Edit, attach file or add comment